OpenStudy (anonymous):

Please answer! Find parametric equations for the curve y = x + 1, x ≥ 0, by using as parameter the distance of the point (x,y) from the origin.

OpenStudy (loser66):

to me, I just do something like x = t y = 1 +t that's it

OpenStudy (anonymous):

That would be the answer?

OpenStudy (loser66):

@e.mccormick

OpenStudy (e.mccormick):

You also need the condition in the parametric answer.

OpenStudy (loser66):

ok, so it t>=0?

OpenStudy (e.mccormick):

Yah, and perhaps \(\in\mathbb{R}\)

OpenStudy (e.mccormick):

Let \(x=t,~ y=t+1;t\in\mathbb{R}\ge 0\)

OpenStudy (e.mccormick):

I think that is a proper answer. Sometimes they want { } around parametic answers as well. Setbuilder notation.

OpenStudy (e.mccormick):

Parametric from Parameter and I have no clue what. Probably just an ending denoting action or an item. Para- beside, -meter measure. So it it the measure of things near together. A set. That is the language of it, which kind of explains why they use this odd term to mean something as simple as "build a set for this with all the rules needed."